# Fluctuations of the Ising free energy on Erdős-Rényi graphs

### v1

@@ 1.1. Background and motivation
In an important contribution Dembo and Montanari [23] proved that to the leading order the free energy of the Ising model on $\mathbb{G}(n,m)$ follows the ‘replica symmetric solution’ predicted by the ‘cavity method’, a heuristic from statistical physics [49, 50]. The replica symmetric solution comes in terms of a fixed point problem on the space $\mathcal{P}([-1,1])$ of probability measures on the interval [1,1][-1,1]. To elaborate, let
@@
in probability. [^1] Actually Dembo and Montanari prove a more general result. Namely, they show that for any sequence of graphs that converges to a unimodular tree in the topology of local weak convergence the free energy converges to the leading order to the replica symmetric solution associated with the limiting unimodular random tree. We will come back to this in Section 2.5. Also the formula (1.5) is a slightly simplified version of the formula from [23]; see Section 2.5 for details.
@@ 1.2.1. The case $B>0$
We prove that the (suitably shifted and scaled) free energy converges to a Gaussian whose variance derives from the fixed points of a family of operators on the space $\mathcal{P}([-1,1]^{2})$ of probability measures on the square $1,1]^{2}$. To elaborate, for a parameter $t\in[0,1]$ define an operator
@@
be sequences of random variables with distribution $\pi$. In addition, let
@@ 1.2.2. The case $B=0$
Indeed, for $\beta<\beta_{\mathrm{f}}(d)$ the distribution $\pi_{d,\beta,0}$ from (1.3) turns out to be just the atom at zero. In effect, the functional (1.4) simply evaluates to
@@
$\lim_{n\to\infty}\log Z_{\mathbb{G}}(\beta,0)-\left[n\log 2+m\log\operatorname{ch}\beta+\frac{1}{2}\log\frac{2}{\pi n}\right]$ $=\sum_{\ell\geq 3}\left({\boldsymbol{Y}_{\ell}\log\left({1+\operatorname{th}^{\ell}\beta}\right)-\frac{d^{\ell}}{2\ell}\operatorname{th}^{\ell}\beta}\right)-\frac{d}{2}\operatorname{th}\beta-\frac{d^{2}}{4}\operatorname{th}^{2}\beta.$ (1.15)
@@
Theorem 1.3 is syntactically similar to Theorem 1.1 but for the appearance of the countable ‘exceptional’ set $\mathfrak{X}_{d}$ of discontinuities of the function $\beta\mapsto r_{d,\beta}$ from [11, 23] where we do not know that convergence to a Gaussian limit occurs. Actually for all we know the exceptional set may be empty, i.e., the function $\beta\mapsto r_{d,\beta}$ may be continuous for all $d>1$; Figure 2 displays $\beta\mapsto r_{d,\beta}$ for a few values of $d$. But of course a proof (or refutation) of this hypothesis remains an open problem.
@@
Additionally, the heat map representation in Figure 3 shows the distributions $\pi_{d,\beta,0,t}$ for three increasing values of $t$. The figure shows how the probability mass gravitates towards the diagonal as $t$ increases. Indeed, as $t$ increases the two coordinates of the random pair
@@
become more ‘aligned’. For instance, for $t=0$ the Poisson variables from (1.8) satisfy $\boldsymbol{d}_{t}^{(0)}=0$ while $\boldsymbol{d}_{t}^{(1)},\boldsymbol{d}_{t}^{(2)}$ are independent ${\rm Po}(d)$ variables. Consequently, comparing the definition (1.9) of the operator $\mathrm{BP}_{d,\beta,B,0}^{\otimes}$ with the definition (1.2) of $\mathrm{BP}_{d,\beta,B}$, we see that the coordinates of the pair (1.18) are just two independent random variables with distribution $\pi_{d,\beta,B}$. On the other extreme, for $t=1$ we have $\boldsymbol{d}_{t}^{(1)}=\boldsymbol{d}_{t}^{(2)}=0$ while $\boldsymbol{d}_{t}^{(0)}$ is a ${\rm Po}(d)$ variable. Hence, the operator (1.9) acts identically on both coordinates, which means that the two coordinates of (1.18) are identical and have distribution $\pi_{d,\beta,B}$. Naturally, for values $0<t<1$ the distribution $\pi_{d,\beta,B,t}^{\otimes}$ interpolates between these two extremes. We make a quick note of these observations for later reference.
@@ 1.3. Notation
We denote the vertex set of the random graph $\mathbb{G}(n,m)$ by $V_{n}=\{v_{1},\ldots,v_{n}\}$. Throughout the paper asymptotic notation such as $o(\,\cdot\,),O(\,\cdot\,)$ refers to the limit $n\to\infty$. The random graph $\mathbb{G}(n,m)$ enjoys a property $\mathfrak{P}$ with high probability (‘w.h.p. ’) if the probability that $\mathfrak{P}$ occurs converges to one as $n\to\infty$. We continue to assume tacitly that $m=m(n)$ is a sequence such that $\lim_{n\to\infty}2m/n=d$ for a fixed real $d>0$.
@@
Throughout the paper we assume tacitly that $\beta>0$ and $B\geq 0$. Additionally, for a graph $G$ rooted at a vertex $r\in V(G)$ and an integer $\ell\geq 0$ we will be led to consider the conditional distribution
$\mu_{G,r,\beta,B}^{(\ell)}(\sigma)$ $=\mu_{G,\beta,B}\left({\sigma\mid\{\forall v\in\partial^{\ell}(G,r):\boldsymbol{\sigma}(v)=1\}}\right)$ $(\sigma\in\{\pm 1\}^{V(G)}\})$ (1.21)
@@ 2. Overview
The basic idea behind the proofs of the central limit theorems Theorems 1.1 and 1.3 is to consider a family of correlated random graphs. The degree of correlation is gauged by a parameter $t\in[0,1]$, with $t=1$ meaning that the two random graphs are identical and $t=0$ corresponding to two independent random graphs. By interpolating from $t=0$ to $t=1$ we will be able to get a handle on the variance of $\log Z_{\mathbb{G}}(\beta,B)$. To obtain the central limit theorems we will ultimately combine the analysis of the variance with an off-the-shelf martingale central limit theorem.
A similar broad proof strategy was recently used to show that the partition function of the random 2-SAT model (i.e., the number of satisfying assignments) has a central limit theorem throughout the regime of clause/variable densities where satisfying assignments likely exist [14]. The fundamental difference between that proof and the present contribution is that the random 2-SAT problem enjoys a very strong spatial mixing property called ‘Gibbs uniqueness’ throughout the satisfiable regime [1]. By contrast, the Ising model on the random graph does not generally have the Gibbs uniqueness property [17, 23]. The most egregious counterexample is the case $B=0$, $d>1$ and $\beta>\beta_{\mathrm{f}}(d)$, i.e., low temperature and no external field. As Proposition 2.15 below demonstrates, in this case two inverse ‘pure states’ (in the sense of [47]) coexist w.h.p., a scenario that amounts to pretty much the opposite of Gibbs uniqueness.
@@ 2.1. First order approximation: the replica symmetric solution
As a preparation for the proofs of Theorems 1.1 and 1.3 we need to appraise ourselves of the workings of the proof of first order formula (1.5) for the free energy and the role that Belief Propagation plays in this proof. As a key step toward (1.5) Dembo and Montanari show that in the case $B>0$ the distribution $\pi_{d,\beta,B}$ from (1.3) is the limit of the empirical distribution of the vertex magnetisations on the random graph. To elaborate, the magnetisation of a vertex $v_{i}$ is defined as
@@
On the first point, let $\mathbb{T}=\mathbb{T}(d)$ be the (possibly infinite) Galton-Watson tree with root $\varrho$ where every vertex begets a ${\rm Po}(d)$ number of children. Moreover, let $\mathbb{T}^{(\ell)}$ be the (finite) tree obtained from $\mathbb{T}$ by deleting all vertices whose distance from $\varrho$ exceeds $\ell$. The following well known lemma states precisely how $\mathbb{T}$ captures the local geometry of $\mathbb{G}$. Recall that $\nabla^{\ell}(\mathbb{G},v)$ is the subgraph of $\mathbb{G}$ induced on the vertices $u$ at distance at most $\ell$ from $v$, rooted at $v$.
@@
The upshot of Lemma 2.3 is that we can work out the root magnetisation $\langle{\boldsymbol{\sigma}(\varrho)},{\mu_{T,\beta,B}^{(\ell)}}\rangle$ of $T$ subject to the all-ones boundary condition ‘bottom up’ from the root magnetisations of the sub-trees $T_{v}$, $v\in\partial(T,\varrho)$. The recurrence (2.4) can be verified directly in a few lines. Alternatively, Lemma 2.3 follows easily from general statements about Belief Propagation [10, 48].
Lemma 2.3 clarifies the combinatorial meaning of the operator $\mathrm{BP}_{d,\beta,B}$ from (1.1). Indeed, due to the recursive nature of the Galton-Watson tree $\mathbb{T}$, the root has a ${\rm Po}(d)$ number of children. Moreover, for each child $v\in\partial(\mathbb{T},\varrho)$ the tree $\mathbb{T}_{v}$ is again nothing but a Galton-Watson tree with offspring distribution ${\rm Po}(d)$. Thus, (1.2) matches the recurrence (2.4) on the random tree $\mathbb{T}$. Consequently, we obtain the following.
@@
In summary, Eq. (2.2) from Proposition 2.1 shows that the empirical distribution $\pi_{\mathbb{G},\beta,B}$ on the random graph converges to the same limit as the root magnetisation of the random tree $\mathbb{T}^{(\ell)}$ that describes the local geometry of $\mathbb{G}$ subject to the all-ones boundary condition at distance $\ell$ from the root. The fact that the all-ones configuration supplies the ‘correct’ starting point is non-trivial and one of the key facts established in [23]. Further, a moment’s reflection shows that for any $d,\beta,B>0$ the limiting distribution $\pi_{d,\beta,B}$ is not just an atom. Rather, the distribution $\pi_{d,\beta,B}$ reflects the inherent randomness of the local geometry of $\mathbb{G}$ around its vertices $v_{i}$. This fact, to which the Ising model on the Erdős-Rényi graph owes much of its mathematical depth, sets the Erdős-Rényi graph apart other graph classes such as integer lattices or random regular graphs [24].
@@
Given replica symmetry and the distribution of the magnetisations, the Bethe free energy (1.4) is generally known to provide a first order approximation to $\log Z_{\mathbb{G}}(\beta,B)$ [20]. This can be verified by a coupling argument called the Aizenman-Sims-Starr scheme [4]. The basic idea is to reduce the calculation of the free energy to the careful computation of the effect on the partition function of a sequence of local changes to the graph, such as the addition of single edges. In combination with knowledge of the distribution of the magnetisations (2.2) the replica symmetry condition (2.6) suffices to compute the impact of such local changes with sufficient precision.
@@ 2.2. The central limit theorem in the case $B>0$
To prove the central limit theorems (Theorems 1.1 and 1.3) we develop a refined version of the Belief Propagation formalism that captures the fluctuations of $\log Z_{\mathbb{G}}(\beta,B)$. In the case $B>0$ we can directly use some of the aforementioned ingredients of the proof of (1.5), particularly Proposition 2.1. However, as we will see in Section 2.3 below, the case $B=0$ turns out to be more delicate. The proofs for the following statements can be found in Section 4, where we prove Lemma 2.3, Proposition 2.5, Lemma 2.7, Lemma 2.8, Proposition 2.9, Lemma 2.11 and Propositions 2.10, 2.12 and 2.13.
We are going to derive Theorem 1.1 from a generic martingale central limit theorem. The main challenge will be to get a handle on the resulting variance process. The martingale that we use is the well known edge exposure martingale [6, Chapter 7]. Hence, we think of the random graph $\mathbb{G}=\mathbb{G}(n,m)$ as being obtained by picking one random edge at a time without replacement. Thus, let $\boldsymbol{e}_{1},\ldots,\boldsymbol{e}_{m}$ be a uniformly random sequence of pairwise distinct edges and let $E(\mathbb{G}(n,m))=\{\boldsymbol{e}_{1},\ldots,\boldsymbol{e}_{m}\}$ be the edge set of the random graph. Furthermore, let $\mathfrak{F}_{n,M}$ be the $\sigma$-algebra generated by the first $M$ edges $\boldsymbol{e}_{1},\ldots,\boldsymbol{e}_{M}$. Hence, given $\mathfrak{F}_{n,M}$ we know $\boldsymbol{e}_{1},\ldots,\boldsymbol{e}_{M}$, while $\boldsymbol{e}_{M+1},\ldots,\boldsymbol{e}_{m}$ remain unexposed, viz. random. Let
@@
Thus, Proposition 2.5 reduces the proof of Theorem 1.1 to the analysis of the variance process $(\boldsymbol{X}_{n,M}^{2})_{1\leq M\leq m}$. To investigate the variance process we adapt a technique originally developed in [14] to study the random 2-SAT problem to the present setting of the Ising model on random graphs. Namely, we will express $\boldsymbol{X}_{n,M}^{2}$ in terms of a family of correlated random graphs. Specifically, we consider three pairs of correlated random graphs
@@
The following three-type Galton-Watson tree $\mathbb{T}^{\otimes}=\mathbb{T}^{\otimes}(d,t)$ fits this bill. The three vertex type are simply type 0, type 1 and type 2. The root $\varrho$ is type 0. A type 0 vertex begets a ${\rm Po}(dt)$ number of type 0 vertices as offspring, as well as ${\rm Po}(d(1-t))$ many type 1 vertices and ${\rm Po}(d(1-t))$ type 2 vertices; these three Poissons are mutually independent. Moreover, for $h=1,2$ the offspring of a type $h$ vertex is just a ${\rm Po}(d)$ number of type $h$ vertices.
We say that two possible outcomes $T,T^{\prime}$ of $\mathbb{T}^{\otimes}$ are isomorphic if there exists a tree isomorphism $T\to T^{\prime}$ that preserves the root and the types. Further, for $h=1,2$ we let $\mathbb{T}^{\otimes}[h]$ be the sub-tree of $\mathbb{T}^{\otimes}$ obtained by deleting all vertices of type 3h3-h. In addition, for $\ell\geq 0$ obtain $\mathbb{T}^{\otimes\,(\ell)}[h]$ from $\mathbb{T}^{\otimes}[h]$ by deleting all vertices whose distance from $\varrho$ exceeds $\ell$.

### v2

@@ 1.1. Background and motivation
In an important contribution Dembo and Montanari [23] proved that to the leading order the free energy of the Ising model on $\mathbb{G}(n,m)$ follows the ‘replica symmetric solution’ predicted by the cavity method, a heuristic from statistical physics [49, 50]. The replica symmetric solution comes in terms of a fixed point problem on the space $\mathcal{P}([-1,1])$ of probability measures on the interval [1,1][-1,1]. To elaborate, let
@@
in probability. [^1] Actually Dembo and Montanari prove a more general result. Namely, they show that for any sequence of graphs that converges to a unimodular tree in the topology of local weak convergence the free energy converges to the leading order to the replica symmetric solution associated with the limiting unimodular random tree. We will come back to this in Section 2.5.
@@ 1.2.1. The case $B>0$
We prove that the (suitably shifted and scaled) free energy converges to a Gaussian whose variance derives from the fixed points of a family of operators on the space $\mathcal{P}([-1,1]^{2})$ of probability measures on the square $1,1]^{2}$. To elaborate, for a parameter $t\in[0,1]$ define the coupled Belief Propagation operator
@@
be sequences of random variables (viz., vectors) with distribution $\pi$. In addition, let
@@ 1.2.2. The case $B=0$
Indeed, for $d\leq 1$ or $\beta<\beta_{\mathrm{f}}(d)$ the distribution $\pi_{d,\beta,0}$ from (1.3) turns out to be just the atom at zero. In effect, the functional (1.4) simply evaluates to
@@
$\lim_{n\to\infty}\log Z_{\mathbb{G}}(\beta,0)-\left[n\log 2+m\log\operatorname{ch}\beta-\frac{1}{2}\log\left({1-d\operatorname{th}\beta}\right)\right]$ $=\sum_{\ell\geq 3}\left({\boldsymbol{Y}_{\ell}\log\left({1+\operatorname{th}^{\ell}\beta}\right)-\frac{d^{\ell}}{2\ell}\operatorname{th}^{\ell}\beta}\right)-\frac{d}{2}\operatorname{th}\beta-\frac{d^{2}}{4}\operatorname{th}^{2}\beta.$ (1.15)
@@
Theorem 1.3 is syntactically similar to Theorem 1.1 but for the appearance of the countable ‘exceptional’ set $\mathfrak{X}_{d}$ of discontinuities of the function $\beta\mapsto r_{d,\beta}$ from (1.16) where we do not know that convergence to a Gaussian limit occurs. Actually for all we know the exceptional set may be empty, i.e., the function $\beta\mapsto r_{d,\beta}$ may be continuous for all $d>1$; Figure 2 displays $\beta\mapsto r_{d,\beta}$ for a few values of $d$. But of course a proof (or refutation) of this hypothesis remains an open problem.
@@
Additionally, the heat map representation in Figure 3 shows the distributions $\pi_{d,\beta,0,t}$ for three increasing values of $t$. The figure shows how the probability mass gravitates towards the diagonal as $t$ increases. Roughly speaking, as $t$ increases the two coordinates of the random pair
@@
become more ‘aligned’. For instance, for $t=0$ the Poisson variables from (1.8) satisfy $\boldsymbol{d}_{t}^{(0)}=0$ while $\boldsymbol{d}_{t}^{(1)},\boldsymbol{d}_{t}^{(2)}$ are independent ${\rm Po}(d)$ variables. Consequently, comparing the definition (1.9) of the operator $\mathrm{BP}_{d,\beta,B,0}^{\otimes}$ with the definition (1.2) of $\mathrm{BP}_{d,\beta,B}$, we see that the coordinates of the pair (1.18) are just two independent random variables with distribution $\pi_{d,\beta,B}$. On the other extreme, for $t=1$ we have $\boldsymbol{d}_{t}^{(1)}=\boldsymbol{d}_{t}^{(2)}=0$ while $\boldsymbol{d}_{t}^{(0)}$ is a ${\rm Po}(d)$ variable. Hence, the operator (1.9) acts identically on both coordinates, which means that the two coordinates of (1.18) are identical and have distribution $\pi_{d,\beta,B}$. Naturally, for values $0<t<1$ the distribution $\pi_{d,\beta,B,t}^{\otimes}$ interpolates between these two extremes. Let us make a quick note of these observations for later reference.
@@ 1.3. Notation
We denote the vertex set of the random graph $\mathbb{G}(n,m)$ by $V_{n}=\{v_{1},\ldots,v_{n}\}$. Throughout the paper asymptotic notation such as $o(\,\cdot\,),O(\,\cdot\,)$ refers to the limit $n\to\infty$. The random graph $\mathbb{G}(n,m)$ enjoys a property $\mathfrak{P}$ with high probability (‘w.h.p. ’) if the probability that $\mathfrak{P}$ occurs converges to one as $n\to\infty$. We continue to assume tacitly that $m=m(n)$ is a sequence such that $\lim_{n\to\infty}2m/n=d$ for a fixed real $d>0$. All statements concerning the random graph are understood to hold for any sequence $m=m(n)$ with $\lim_{n\to\infty}2m/n=d$.
@@
Additionally, for a graph $G$ rooted at a vertex $r\in V(G)$ and an integer $\ell\geq 0$ we will be led to consider the conditional distribution
$\mu_{G,r,\beta,B}^{(\ell)}(\sigma)$ $=\mu_{G,\beta,B}\left({\sigma\mid\{\forall v\in\partial^{\ell}(G,r):\boldsymbol{\sigma}(v)=1\}}\right)$ $(\sigma\in\{\pm 1\}^{V(G)})$ (1.21)
@@
From here on we assume tacitly that $d,\beta>0$ and $B\geq 0$. Unless specified otherwise all statements are understood to hold for any such $d,\beta,B$.
@@ 2. Overview
The basic idea behind the proofs of the central limit theorems Theorems 1.1 and 1.3 is to consider a family of correlated random graphs. The degree of correlation is gauged by a coupling parameter $t\in[0,1]$, with $t=1$ meaning that the two random graphs are identical and $t=0$ corresponding to two independent random graphs. By interpolating from $t=0$ to $t=1$ we will be able to get a handle on the variance of $\log Z_{\mathbb{G}}(\beta,B)$. To obtain the central limit theorems we will ultimately combine the analysis of the variance with an off-the-shelf martingale central limit theorem.
A broadly similar proof strategy was recently used to show that the partition function of the random 2-SAT model (i.e., the number of satisfying assignments) has a central limit theorem throughout the entire regime of clause/variable densities where satisfying assignments likely exist [14]. The fundamental difference between that proof and the present contribution is that the random 2-SAT problem enjoys a very strong spatial mixing property called ‘Gibbs uniqueness’ throughout the satisfiable regime [1]. The Gibbs uniqueness property constitutes the linchpin of the proofs from [14]. By contrast, the Ising model on the random graph does not have the Gibbs uniqueness property for most interesting parameter regimes [17, 23]. The most egregious counterexample is the case $B=0$, $d>1$ and $\beta>\beta_{\mathrm{f}}(d)$, i.e., low temperature and no external field. As Proposition 2.15 below demonstrates, in this case two inverse ‘pure states’ (in the sense of [47]) coexist w.h.p., a scenario that amounts to pretty much the opposite of Gibbs uniqueness. Additionally, the Belief Propagation operator has a spurious (and unstable) fixed point at $\delta_{0}$.
@@ 2.1. First order approximation: the replica symmetric solution
As a preparation for the proofs of Theorems 1.1 and 1.3 we need to appraise ourselves of the workings of the proof of first order formula (1.5) for the free energy and the role that Belief Propagation plays in this proof. As a key step toward (1.5) Dembo and Montanari show that in the case $B>0$ the distribution $\pi_{d,\beta,B}$ from (1.3) is the limit of the empirical distribution of the vertex magnetisations on the random graph. To elaborate, the magnetisation of a vertex $v_{i}$ in $\mathbb{G}$ is defined as
@@
On the first point, let $\mathbb{T}=\mathbb{T}(d)$ be the (possibly infinite) Galton-Watson tree with root $\varrho$ where every vertex begets a ${\rm Po}(d)$ number of children. Moreover, let $\mathbb{T}^{(\ell)}$ be the (finite) tree obtained from $\mathbb{T}$ by deleting all vertices whose distance from $\varrho$ exceeds $\ell$. The following well-known lemma states precisely how $\mathbb{T}$ captures the local geometry of $\mathbb{G}$. Recall that $\nabla^{\ell}(\mathbb{G},v)$ is the subgraph of $\mathbb{G}$ induced on the vertices $u$ at distance at most $\ell$ from $v$, rooted at $v$.
@@
The upshot of Lemma 2.3 is that we can work out the root magnetisation $\langle{\boldsymbol{\sigma}(\varrho)},{\mu_{T,\beta,B}^{(\ell)}}\rangle$ of $T$ subject to the all-ones boundary condition ‘bottom up’ from the root magnetisations of the sub-trees $T_{v}$, $v\in\partial(T,\varrho)$. The recurrence (2.4) can be verified directly in a few lines (see Section 4.1). Alternatively, Lemma 2.3 could be derived easily from general statements about Belief Propagation [10, 48].
Lemma 2.3 clarifies the combinatorial meaning of the operator $\mathrm{BP}_{d,\beta,B}$ from (1.1). Indeed, due to the recursive nature of the Galton-Watson tree $\mathbb{T}$, the root has a ${\rm Po}(d)$ number of children. Moreover, for each child $v\in\partial(\mathbb{T},\varrho)$ the tree $\mathbb{T}_{v}$ is again nothing but a Galton-Watson tree with offspring distribution ${\rm Po}(d)$. Thus, (1.2) matches the recurrence (2.4) on the random tree $\mathbb{T}$. Consequently, we obtain the following immediate consequence of Lemma 2.3.
@@
In summary, Eq. (2.2) from Proposition 2.1 shows that the empirical distribution $\pi_{\mathbb{G},\beta,B}$ on the random graph converges to the same limit as the root magnetisation of the random tree $\mathbb{T}^{(\ell)}$ that describes the local geometry of $\mathbb{G}$ subject to the all-ones boundary condition at distance $\ell$ from the root. The fact that the all-ones configuration supplies the ‘correct’ starting point is non-trivial and one of the key facts established in [23]. Further, a moment’s reflection shows that for $B>0$ the limiting distribution $\pi_{d,\beta,B}$ is not just an atom. Rather, the distribution $\pi_{d,\beta,B}$ reflects the inherent randomness of the local geometry of $\mathbb{G}$ around its vertices $v_{i}$. This fact, to which the Ising model on the Erdős-Rényi graph owes much of its mathematical depth, sets the Erdős-Rényi graph apart from other graph classes such as integer lattices or random regular graphs [24].
@@
Given replica symmetry and the distribution of the magnetisations, the Bethe free energy (1.4) is generally known to provide a first order approximation to $\log Z_{\mathbb{G}}(\beta,B)$ [20]. This can be verified by a coupling argument called the Aizenman-Sims-Starr scheme [4]. The basic idea behind this technique is to reduce the calculation of the free energy to the careful computation of the effect on the partition function of a sequence of local changes to the graph, such as additions of single edges. In combination with knowledge of the distribution of the magnetisations (2.2) the replica symmetry condition (2.6) suffices to compute the impact of such local changes accurately.
@@ 2.2. The central limit theorem in the case $B>0$
To prove the central limit theorems (Theorems 1.1 and 1.3) we develop a refined version of the Belief Propagation formalism that captures the fluctuations of $\log Z_{\mathbb{G}}(\beta,B)$. In the case $B>0$ we can directly use some of the aforementioned ingredients of the proof of (1.5), particularly Proposition 2.1. However, as we will see in Section 2.3 below, the case $B=0$ turns out to be more delicate. The proofs for the following statements can be found in Section 4, where we prove Lemma 2.3, Proposition 2.5, Lemma 2.7, Lemma 2.8, Proposition 2.9, Lemma 2.11 and Propositions 2.10, 2.12 and 2.13. Some of the intermediate statements in this subsection actually also hold for $B=0$, in which case we will be able to reuse them in Section 2.3 towards the proof of Theorem 1.3.
We are going to derive Theorem 1.1 from a generic martingale central limit theorem. The main challenge will be to get a handle on the resulting variance process. The martingale that we use is the well-known edge exposure martingale [6, Chapter 7]. Hence, we think of the random graph $\mathbb{G}=\mathbb{G}(n,m)$ as being obtained by picking one random edge at a time without replacement. Thus, let $\boldsymbol{e}_{1},\ldots,\boldsymbol{e}_{m}$ be a uniformly random sequence of pairwise distinct edges and let $E(\mathbb{G}(n,m))=\{\boldsymbol{e}_{1},\ldots,\boldsymbol{e}_{m}\}$ be the edge set of the random graph. Furthermore, let $\mathfrak{F}_{n,M}$ be the $\sigma$-algebra generated by the first $M$ edges $\boldsymbol{e}_{1},\ldots,\boldsymbol{e}_{M}$. Hence, given $\mathfrak{F}_{n,M}$ we know $\boldsymbol{e}_{1},\ldots,\boldsymbol{e}_{M}$, while $\boldsymbol{e}_{M+1},\ldots,\boldsymbol{e}_{m}$ remain unexposed, viz. random (albeit not quite independent of $\boldsymbol{e}_{1},\ldots,\boldsymbol{e}_{M}$ because edges are chosen without replacement). Let
@@
Thus, Proposition 2.5 reduces the proof of Theorem 1.1 to the analysis of the variance process $(\boldsymbol{X}_{n,M}^{2})_{1\leq M\leq m}$. To investigate the variance process we adapt a technique originally developed in [14] to study the random 2-SAT problem to the present setting of the Ising model on random graphs. Namely, we will express $\boldsymbol{X}_{n,M}^{2}$ in terms of a family of coupled random graphs. Specifically, we consider three pairs of random graphs
@@
The following three-type Galton-Watson tree $\mathbb{T}^{\otimes}=\mathbb{T}^{\otimes}(d,t)$ fits this bill. The three vertex types are simply type 0, type 1 and type 2. The root $\varrho$ is type 0. A type 0 vertex begets a ${\rm Po}(dt)$ number of type 0 vertices as offspring, as well as ${\rm Po}(d(1-t))$ many type 1 vertices and ${\rm Po}(d(1-t))$ type 2 vertices; these three Poissons are mutually independent. Moreover, for $h=1,2$ the offspring of a type $h$ vertex is just a ${\rm Po}(d)$ number of type $h$ vertices.
We say that two possible outcomes $T,T^{\prime}$ of $\mathbb{T}^{\otimes}$ are isomorphic if there exists a tree isomorphism $T\to T^{\prime}$ that preserves the root and the types. Further, for $h=1,2$ we let $\mathbb{T}^{\otimes}[h]$ be the sub-tree of $\mathbb{T}^{\otimes}$ obtained by deleting all vertices of type 3h3-h. In addition, for $\ell\geq 0$ we obtain $\mathbb{T}^{\otimes\,(\ell)}[h]$ from $\mathbb{T}^{\otimes}[h]$ by deleting all vertices whose distance from $\varrho$ exceeds $\ell$.